A 224 kg crate is pushed horizontally with a force of 710 N. If the coefficient of friction is 0.25, calculate the acceleration of the crate.
1
Expert's answer
2010-11-19T06:04:58-0500
According the Second Newton's law
ma = F - μ mg;
a = F/m - μ g = 710/224 - 0.25* 10 = 0.67 m/s2